OpenStudy (anonymous):

use the second equation to make a substitution for y in the first equation. 2x + y = 6 y = 3x + 4 What is the resulting equation?

OpenStudy (anonymous):

one sec i know who can help you out.

OpenStudy (anonymous):

@shyguybros

OpenStudy (anonymous):

@Melodysim

OpenStudy (anonymous):

thank you :) i appreciate it

OpenStudy (anonymous):

Welcome!!!!

OpenStudy (anonymous):

Okay, so you're going to want to get "x" by itself

OpenStudy (anonymous):

Er wait

OpenStudy (anonymous):

Sorry, I'm not sure how to do this

OpenStudy (anonymous):

its okey dont worry.

OpenStudy (jessicajohnson):

2x + (3x + 4) = 6 5x +4 =6 This would be the resulting equation after substitution 5x = 2 x = 2/5 but I went ahead and solved it

OpenStudy (anonymous):

thank you so much jessica :)

OpenStudy (jessicajohnson):

no problem!

OpenStudy (anonymous):

Thank you Jessica, now I remember how to do this :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I completely forgot about that haha

OpenStudy (jessicajohnson):

:)